OpenStudy (anonymous):

Solve v-(v+4)+2v=v-7-v+19

OpenStudy (anonymous):

v-v-4+2v=v-7-v+19 -4+2v=12 2v=16 v=8

OpenStudy (anonymous):

Thanks I was confused on the variables but I know what to do now so thanks again.